# Supplemental Material: Learning Subject-Aware Cropping by Outpainting Professional Photos

### v1

@@ A Additional Results and Ablations
Table 4: Comparison on the 50 human-centric test images in GAICD. * indicates results reported by (Zhang et al. 2022); refer to their paper for the full table of baselines. As noted in supplemental §A.9, the sparse supervision generated by GenCrop is poorly calibrated to the ranking labels in GAICD. Our version of GenCrop (GenCrop-R), modified for ranking performs better than VFN, which also does not have access to score and rank labels (e.g., being trained on Flickr or Unsplash), but results are poor compared to models that directly train on GAICD.
@@
We also provide examples of generated images from our pipeline. Figs. 56, and 7 show outpainted images produced by our pipeline and used for training; images outpainted without text-conditioning; and images that are discarded by filtering, respectively. We hope that these additional results will provide useful baselines and commentary for future research on weakly-supervised image cropping.
@@ A.9 Comparison on Crop Ranking Tasks
Figure 6: Examples of outpainting with (left) and without (right) text conditioning. Conditioning with an estimated caption from BLIP-2 (Li et al. 2023) improves the quality of the synthesized region. Without text conditioning, Stable Diffusion (Rombach et al. 2022) may halucinate arbitary other objects and people into the oupainted regions. The text can be approximate since its purpose is to prevent arbitrary content in the halucinated region and the subject region is already known.

### v2

# Learning Subject-Aware Cropping by Outpainting Professional Photos

@@ Appendix A Additional Results and Ablations
Table 9: Comparison on the 50 human-centric test images in GAICD. * indicates results reported by (Zhang et al. 2022); refer to their paper for the full table of baselines. As noted in §A.9, the sparse supervision generated by GenCrop is poorly calibrated to the ranking labels in GAICD. Our version of GenCrop (GenCrop-R), modified for ranking performs better than VFN, which also does not have access to score and rank labels (e.g., being trained on Flickr or Unsplash), but results are poor compared to models that directly train on GAICD.
@@
We also provide examples of generated images from our pipeline. Figs. 1112, and 13 show outpainted images produced by our pipeline and used for training; images outpainted without text-conditioning; and images that are discarded by filtering, respectively. We hope that these additional results will provide useful baselines and commentary for future research on weakly-supervised image cropping.
@@ A.9 Comparison on Crop Ranking Tasks
Figure 12: Examples of outpainting with (left) and without (right) text conditioning. Conditioning with an estimated caption from BLIP-2 (Li et al. 2023) improves the quality of the synthesized region. Without text conditioning, Stable Diffusion (Rombach et al. 2022) may hallucinate arbitrary other objects and people into the outpainted regions. The text can be approximate since its purpose is to prevent arbitrary content in the hallucinated region and the subject region is already known.